OpenStudy (anonymous):

2/12 - (-3/4) = ? (Show Work)

OpenStudy (sasogeek):

do you have an idea how to go about solving this?

OpenStudy (anonymous):

no...

OpenStudy (sasogeek):

you have asked several of these in a row, i believe those who helped you have shown you HOW to go about solving this kind of problem... haven't they?

OpenStudy (anonymous):

I am checking my work on here

OpenStudy (sasogeek):

what is your answer? I'll let you know if it's correct since you're only checking :)

OpenStudy (anonymous):

never mind my brother says I did mess up....I HATE MATH! -3-

OpenStudy (anonymous):

2/12-(-3/4) =4/24-(-18/24) =4/24+18/24 =22/24 =11/12

OpenStudy (anonymous):

DAME I WAS WAY OFF!

OpenStudy (sasogeek):

person, just a caution, you are not supposed to give out answers to users... usually you're supposed to guide them to understanding how to arrive at answers, not giving it the answers just as they are :) be sure to remember that

OpenStudy (anonymous):

oops, sorry

OpenStudy (anonymous):

thanks guy for helping!

OpenStudy (sasogeek):

it's okay, that's why it is just a caution for now :)

OpenStudy (sasogeek):

do you understand how to solve such questions now? kelly...

OpenStudy (anonymous):

yah...i think

OpenStudy (sasogeek):

ok, quick question. tell me what you get. (5/3) - (4/5)

OpenStudy (anonymous):

hold up let me check my notes

OpenStudy (anonymous):

hmmmm.....

OpenStudy (sasogeek):

..........?